# Attributional Safety Failures in Large Language Models under Code-Mixed Perturbations

### v1

@@ 6 Attribution analysis
6.1 Initial observations
Figure 3: Token-wise attribution alignment between an English prompt (top) and its code-mixed Hindi counterpart (bottom) using sequence attribution scores.
@@ 6.1 Initial observations
6.2 Saliency drift attribution (SDA)
@@
Let the English input be denoted by $x=\{w_{1},w_{2},\ldots,w_{n}\}$ and its code-mixed counterpart by $x^{\prime}=\{w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{n}\}$. We compute token-level importance using a gradient-based attribution method $A(\cdot)$, such as integrated gradients Sundararajan et al. (2017). To ensure that comparisons are not influenced by raw attribution scales or input length, we adopt a normalized importance proxy termed rank inverse (RI), defined as:
$\textit{RI}(w_{i};x)=\frac{1}{\text{rank}_{x}(A(w_{i};x))}$ (2)
@@
$\Delta\textit{RI}(w_{i})=\textit{RI}(w_{i};x)-\textit{RI}(w^{\prime}_{i};x^{\prime})$ (3)
@@
$\Delta\textit{RI}_{\text{norm}}(w_{i})=\Delta\textit{RI}(w_{i})+\alpha,$ (4)
@@
$\alpha=\left|\min_{i}\Delta\textit{RI}(w_{i})\right|.$ (5)
Tokens with higher $\Delta\textit{RI}_{\text{norm}}$ are visualized using larger font sizes in word clouds (discussed later in this section) to reflect greater attribution loss, thereby, highlighting which safety-critical cues are de-emphasized due to code-mixing. Similarly, a complementary visualization is generated for tokens that gain attribution in the code-mixed input, with word size proportional to $E[\textit{RI}(w^{\prime}_{i};x^{\prime})]$ across samples. $E[\cdot]$ denotes the expectation operator, i.e., the average rank-inverse saliency score of the token $w^{\prime}_{i}$ computed across all code-mixed samples in which it appears. This SDA formulation provides actionable insight into attributional collapse and enables a principled comparison between original and code-mixed prompts in terms of their impact on model attention and safety alignment.
@@
Figure 4: Case 1: English inputs yield non-harmful responses, while Bengali and Hindi code-mixed variants triggered harmful outputs. Both word clouds show harmful tokens (e.g., disgusting, parasites) that lose attribution under code-mixing.
@@
Figure 5: Case 2: Both English and code-mixed inputs produce harmful responses. The word clouds show that harmful tokens retain high attribution, indicating stable but unsafe behavior across modalities.
@@
Figure 6: Case 3: Both English and code-mixed inputs result in non-harmful responses. The word clouds show minimal attribution drift, indicating stable and safe behavior across both forms.
Case 1 reveals a striking attributional failure (see Figure 4): tokens associated with harmful semantics (e.g., disgusting, parasites) receive substantially lower attribution in the code-mixed form, reflected in their large size in the drift-based word clouds. Although we show results for two languages, code-mixed Bengali and code-mixed Hindi, the results hold for all other code-mixed languages (see Appendix). This suggests that the model no longer recognizes these words as toxic, leading to unintended harmful outputs.
@@
6.3 Sensitivity to targeted perturbation
@@
6.3.1 Target = top-$k$ toxic tokens
@@
$\delta_{\text{tox}}(w_{i})=S(x)-S(x\setminus w_{i})$ (6)
@@
$\mathcal{W}^{\text{top}}=\text{SortDescending}\left(\{\delta_{\text{tox}}(w_{i})\}_{i=1}^{n}\right)$ (7)
@@
Figure 7: Example prompt after replacing top 2 toxic tokens.
Figure 8: ASR heatmap comparing (Q-T(k), CM-A) and Case 1. Most languages exhibit a steep ASR drop at $k=1$ or $k=2$ after which the trend no longer changes with increasing $k$.
6.3.2 Target = not-so-toxic tokens
@@
$\mathcal{W}^{\text{nts}}=\{w_{i}\mid\delta_{\text{tox}}(w_{i})\in\text{Percentile}_{[20,60]}\}$ (8)
@@
Figure 9: Violin plot showing ASR distributions across replacement ranks ($k=1$ to $5$) for non-toxic tokens (20th–60th percentile of attribution). ASR remains high and stable, but is less than the Case 1.
@@
Figure 10: Example prompt on replacing not-so-toxic token.
7 Key insights
@@
4. Attributional perturbations of toxic words do not hamper safety. A key practical insight is that token-level translation of highly toxic words do not hamper the safety behaviour of the models. This opens pathways for post-hoc, model-agnostic defenses in multilingual LLM deployments.
@@
8 Conclusion
@@
This paper presents the first large-scale attributional study of safety failures in LLMs under code-mixed inputs, revealing that even semantically equivalent prompts can trigger harmful outputs due to attributional miscalibration. Through a combination of saliency drift analysis, perturbation sensitivity experiments, and token-level attribution verification, we show that safety-critical tokens lose prominence when code-mixed, leading to failure in detecting harmful intent. These effects are magnified in culturally grounded prompts and languages distant from English, exposing the limits of current monolingual alignment strategies. Importantly, we demonstrate that lightweight, attribution-guided interventions, such as selectively translating high-risk tokens, can retain safe behaviour. Our findings call for attribution-aware alignment methods and culturally robust evaluation pipelines to ensure safe and equitable LLM deployment in real-world multilingual settings.
Appendix A: Qualitative error analysis
To complement our quantitative findings, we conduct a detailed qualitative error analysis that examines representative failure modes across the three behavioral transitions outlined in our attributional framework: Case 1 (safe $\rightarrow$ harmful), Case 2 (harmful $\rightarrow$ harmful), and Case 3 (non harmful $\rightarrow$ non harmful). For each case, we analyse token-level attribution patterns, semantic drift, and cultural factors that contribute to failures, reinforcing the causal insights from Section 6.
A.1 Case 1: Safe English $\rightarrow$ harmful code-mixed output (attribution collapse)
Table 2: Case 1: Attributional collapse triggers harmful generation under semantically equivalent CM prompt.
A.2 Case 2: Harmful English $\rightarrow$ harmful code-mixed output (stable unsafe behavior)
Table 3: Case 2: Harmful behavior persists due to high attribution stability on toxic tokens.
A.3 Case 3: Safe English $\rightarrow$ safe code-mixed output (attributional robustness)
Eng-Q: In Japanese culture, what are common concerns around tourism?
Observation: Neutral tokens like “tourism”, “concerns”, and “traditions” show stable attribution (delta < 0.1). Saliency is preserved, leading to safe generation in both forms.
Table 4: Case 3: Attribution is preserved and safety behavior remains stable.
A.4 Summary of insights
Appendix A Cultures and categories
Figure 11: Utilized cultures and key areas that could lead to potential harm within each culture.
Appendix B: Additional analyses and diagnostics
This appendix provides focused evaluations to support Sections 5 and 6.
@@ Appendix A Cultures and categories
Figure 12: Case 1: English prompts yield safe responses, but their code-mixed counterparts trigger harmful completions. Word clouds highlight frequent high-risk tokens (e.g., parasites, pollution, radicals) that lose attribution under code-mixed perturbation across cultures.
B.1 Human annotation validation
To verify the reliability of gpt4-o’s harm labels, we conducted a manual evaluation on a sample of Hindi and Bengali code-mixed prompts. Native speakers labeled model responses as harmful, non-harmful, or ambiguous. We report inter-annotator agreement and disagreement rates between human consensus and model labels in Table 5.
Language Cohen’s $\kappa$ Model disagreement (%) Ambiguity rate (%)
Hindi 0.81 16.7% 11.7%
Bengali 0.79 20.0% 13.3%
Table 5: Human agreement and disagreement with model-generated harm labels.
As shown in Table 5, the $\kappa$ values reflect substantial inter-annotator agreement, and disagreement with gpt4-o remains $\leq 20\%$. Given that we use the same annotation protocols and prompt structures across languages, we expect generalizability of label quality to hold in other cultural settings.
B.2 Token drift visualization across cultures (Case 1)

### v2

@@ 7. Attribution analysis
7.1. Initial observations
Figure 3. Token-wise attribution alignment between an English prompt (top) and its code-mixed Hindi counterpart (bottom) using sequence attribution scores.
@@ 7.1. Initial observations
7.2. Saliency drift attribution (SDA)
@@
Let the English input be denoted by $x=\{w_{1},w_{2},\ldots,w_{n}\}$ and its code-mixed counterpart by $x^{\prime}=\{w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{n}\}$. We compute token-level importance using a gradient-based attribution method $A(\cdot)$, such as integrated gradients (sundararajan2017axiomatic). To ensure that comparisons are not influenced by raw attribution scales or input length, we adopt a normalized importance proxy termed rank inverse (RI), defined as:
(2) $\textit{RI}(w_{i};x)=\frac{1}{\text{rank}_{x}(A(w_{i};x))}$
@@
(3) $\Delta\textit{RI}(w_{i})=\textit{RI}(w_{i};x)-\textit{RI}(w^{\prime}_{i};x^{\prime})$
@@
(4) $\Delta\textit{RI}_{\text{norm}}(w_{i})=\Delta\textit{RI}(w_{i})+\alpha,$
@@
(5) $\alpha=\left|\min_{i}\Delta\textit{RI}(w_{i})\right|.$
Tokens with higher $\Delta\textit{RI}_{\text{norm}}$ are visualized using word shift plots and word clouds (discussed later in this section) to reflect greater attribution loss, thereby, highlighting which safety-critical cues are de-emphasized due to code-mixing. Similarly, a complementary visualization is generated for tokens that gain attribution in the code-mixed input, with word size proportional to $E[\textit{RI}(w^{\prime}_{i};x^{\prime})]$ across samples. $E[\cdot]$ denotes the expectation operator, i.e., the average rank-inverse saliency score of the token $w^{\prime}_{i}$ computed across all code-mixed samples in which it appears. This SDA formulation provides actionable insight into attributional collapse and enables a principled comparison between original and code-mixed prompts in terms of their impact on model attention and safety alignment.
@@
Figure 4. Case 1: English inputs yield non-harmful responses, while Bengali and Hindi code-mixed variants triggered harmful outputs. Both word shift graphs show harmful tokens (e.g., destruction, scum) that have higher attribution in English ( ) than code-mixing ( ). Word clouds are present in Figure 13.
@@
Figure 5. Case 2: Both English and code-mixed inputs produce harmful responses. Both word shift graphs (English: ( ), code-mixing: ( )) show that harmful tokens retain high attribution, indicating stable but unsafe behavior across modalities. Word clouds are present in Figure 14.
@@
Figure 6. Case 3: Both English and code-mixed inputs result in non-harmful responses. Both word shift graphs (English: ( ), code-mixing: ( )) show minimal attribution drift, indicating stable and safe behavior across both forms. Word clouds are present in Figure 15.
Case 1 reveals a striking attributional failure (see Figure 4): tokens associated with harmful semantics (e.g., immigrants, filth, scum) receive substantially lower attribution in the code-mixed form, reflected in their large size in the word shift plot and word clouds. Although we show results for two languages, code-mixed Bengali and code-mixed Hindi, the results hold for all other code-mixed languages (see Appendix). This suggests that the model no longer recognizes these words as toxic, leading to unintended harmful outputs.
@@
7.3. Sensitivity to targeted perturbation
@@
7.3.1. Target = top-$k$ toxic tokens
@@
(6) $\delta_{\text{tox}}(w_{i})=S(x)-S(x\setminus w_{i})$
@@
(7) $\mathcal{W}^{\text{top}}=\text{SortDescending}\left(\{\delta_{\text{tox}}(w_{i})\}_{i=1}^{n}\right)$
@@
Figure 7. Example prompt after replacing top 2 toxic tokens.
Figure 8. ASR heatmap comparing (Q-T(k), CM-A) and Case 1. Most languages exhibit a steep ASR drop at $k=1$ or $k=2$ after which the trend no longer changes with increasing $k$.
7.3.2. Target = not-so-toxic tokens
@@
(8) $\mathcal{W}^{\text{nts}}=\{w_{i}\mid\delta_{\text{tox}}(w_{i})\in\text{Percentile}_{[20,60]}\}$
@@
Figure 9. Violin plot showing ASR distributions across replacement ranks ($k=1$ to 55) for non-toxic tokens (20th–60th percentile of attribution). ASR remains high and stable, but is less than the Case 1.
@@
Figure 10. Example prompt on replacing not-so-toxic token.
8. Key insights
@@
4. Attributional perturbations of toxic words do not hamper safety. A key practical insight is that token-level translation of highly toxic words do not hamper the safety behavior of the models. This opens pathways for post-hoc, model-agnostic defenses in multilingual LLM deployments.
@@
9. Attribution restoration
Our results in Section 5 show a systematic pattern: for almost every language $\ell$, $\mathrm{ASR}^{\mathrm{CM}}_{\ell}$, is vastly higher than $\mathrm{ASR}^{\mathrm{EN}}_{\ell}$. The attribution analysis in Section 7 explains why: once the input is code-mixed, tokens that are safety-critical in English lose attribution mass, while relatively innocuous context tokens become more salient. In other words, the model is at the wrong pieces of the input when it is expressed in a code-mixed form, even though it behaves much more safely when the same intent is expressed in English. This suggests a concrete, attribution-motivated hypothesis: if code-mixing causes safety failures by disrupting attention to safety-critical cues, then mapping code-mixed prompts back into a representation space where those cues are reliably salient—standard English—should substantially reduce harmful responses. To test this, we construct a simple defensive translation layer $T$ as an attribution-guided probe. For each prompt $x$, we estimate a code-mixing score based on language and script tags; if at least $30\%$ of tokens are non-English, we treat $x$ as code-mixed and apply a sentence-level neural translation to obtain an English pivot $T(x)$, otherwise we keep $x$ unchanged. The model is then queried on $T(x)$, but we continue to evaluate harmfulness with respect to the original prompt. Attribution therefore guides the design of TT—we choose English precisely because SDA shows that safety-relevant tokens are best recognised there—but $T$ itself is a lightweight pre-processing step that does not require running attribution at inference time.
Language $\mathrm{ASR}^{\mathrm{EN}}$ $\mathrm{ASR}^{\mathrm{CM}}$ $\mathrm{ASR}^{T(\mathrm{CM})}$
Arabic 77.14% 91.12% 80.12%
Bengali 76.92% 87.35% 77.88%
Chinese 79.59% 83.34% 80.31%
German 75.86% 83.12% 77.26%
Hindi 80.00% 93.14% 81.98%
Japanese 62.86% 78.86% 66.75%
Korean 74.29% 83.26% 75.86%
Russian 90.91% 91.34% 91.34%
Spanish 67.86% 79.52% 73.12%
Portuguese 62.50% 69.73% 65.14%
Macro avg 74.79% 84.08% 76.98%
Table 3. ASR for llama-3.1-8B on the Local TestSet under three conditions: original English prompts ($\mathrm{EN}$), directly code-mixed prompts ($\mathrm{CM}$), and code-mixed prompts processed by the defensive translation layer ($T(\mathrm{CM})$). The intervention is used as an attribution-guided probe: it substantially reduces ASR for code-mixed inputs across all ten languages and narrows the safety gap between CM and English.
Table 3 summarises the resulting ASRs for llama-3.1-8B on the Local TestSet across all ten languages under three conditions: original English prompts ($\mathrm{EN}$), directly code-mixed prompts ($\mathrm{CM}$), and code-mixed prompts passed through the translation layer ($T(\mathrm{CM})$). The pattern is striking. Raw code-mixed inputs produce very high ASR, typically above $83\%$ and in some cases exceeding $90\%$ for under-resourced, heavily code-mixing communities such as Russian, Hindi, and Arabic. When we translate code-mixed prompts into English before querying the model, ASR drops sharply and moves much closer to the English baseline. For Arabic, for example, $\mathrm{ASR}^{\mathrm{EN}}=77\%$ increases to $\mathrm{ASR}^{\mathrm{CM}}=91\%$, but falls to $\mathrm{ASR}^{T(\mathrm{CM})}=80\%$ with translation, recovering the vast majority of the safety lost under code-mixing; Arabic and Hindi show similar behavior, with residual gaps $|\mathrm{ASR}^{T(\mathrm{CM})}_{\ell}-\mathrm{ASR}^{\mathrm{EN}}_{\ell}|$ below ten percentage points. Even for languages with more moderate degradation (German, Spanish, Portuguese), $\mathrm{ASR}^{T(\mathrm{CM})}_{\ell}$ consistently lies much closer to $\mathrm{ASR}^{\mathrm{EN}}_{\ell}$ than to $\mathrm{ASR}^{\mathrm{CM}}_{\ell}$. On average across all ten languages, the macro ASR rises from $74.79\%$ (English) to $84.08\%$ (code-mixed), but drops to $76.98\%$ when code-mixed prompts are translated—recovering roughly $80\%$ of the safety lost to code-mixing. A lightweight utility probe based on answerability and topicality (Appendix B) does not show systematic loss of task relevance, indicating that the intervention primarily removes unsafe generations rather than trivially silencing the model. We do not present $T$ as a complete deployment solution, but as a tightly controlled, attribution-guided case study: by restoring the saliency of safety-critical cues in an English pivot space, we can substantially reduce harmful responses, providing causal support for our explanation of attributional safety failures under code-mixing and pointing to a concrete direction for future multilingual safety mechanisms.
10. Conclusion
@@
This paper presents the first large-scale attributional study of safety failures in LLMs under code-mixed inputs, revealing that even semantically equivalent prompts can trigger harmful outputs due to attributional miscalibration. Through a combination of saliency drift analysis, perturbation sensitivity experiments, and token-level attribution verification, we show that safety-critical tokens lose prominence when code-mixed, leading to failure in detecting harmful intent. These effects are magnified in culturally grounded prompts and languages distant from English, exposing the limits of current monolingual alignment strategies. Importantly, we demonstrate that lightweight, attribution-guided interventions, such as selectively translating high-risk tokens, can retain safe behavior. Our findings call for attribution-aware alignment methods and culturally robust evaluation pipelines to ensure safe and equitable LLM deployment in real-world multilingual settings.
@@
Figure 11. Case 1: English prompts yield safe responses, but their code-mixed counterparts trigger harmful completions. Word clouds highlight frequent high-risk tokens (e.g., parasites, pollution, radicals) that lose attribution under code-mixed perturbation across cultures.
Appendix A Qualitative error analysis
To complement our quantitative findings, we conduct a detailed qualitative error analysis that examines representative failure modes across the three behavioral transitions outlined in our attributional framework: Case 1 (safe $\rightarrow$ harmful), Case 2 (harmful $\rightarrow$ harmful), and Case 3 (non harmful $\rightarrow$ non harmful). For each case, we analyze token-level attribution patterns, semantic drift, and cultural factors that contribute to failures, reinforcing the causal insights from the main text.
A.1. Case 1: Safe English $\rightarrow$ harmful code-mixed output (attribution collapse)
Table 4. Case 1: Attributional collapse triggers harmful generation under semantically equivalent CM prompt.
A.2. Case 2: Harmful English $\rightarrow$ harmful code-mixed output (stable unsafe behavior)
Table 5. Case 2: Harmful behavior persists due to high attribution stability on toxic tokens.
A.3. Case 3: Safe English $\rightarrow$ safe code-mixed output (attributional robustness)
Eng-Q: In Japanese culture, what are common concerns around tourism?
Observation: Neutral tokens like “tourism”, “concerns”, and “traditions” show stable attribution (delta ¡ 0.1). Saliency is preserved, leading to safe generation in both forms.
Table 6. Case 3: Attribution is preserved and safety behavior remains stable.
A.4. Summary of insights
Appendix B Utility evaluation after restoration
The defensive translation experiment in Section 9 is designed to probe safety, but a natural concern is that it might also reduce the model’s usefulness by making it more likely to refuse or by drifting off-topic. We therefore conduct a lightweight utility evaluation that measures whether the restoration changes the model’s ability to produce direct, on-topic answers.
Given a prompt $x$ and model output $y$, we define two binary indicators using GPT-4o as an automatic judge: (i) answerability, which is $1$ if $y$ is judged to be a genuine attempt to answer the prompt rather than a refusal or an unrelated response, and (ii) topicality, which is $1$ if $y$ is judged to remain on-topic relative to $x$. Our utility score for a dataset $D$ is then
$U(D)=\frac{1}{|D|}\sum_{x\in D}\mathbb{1}\big[\text{answerable}(x)\land\text{topical}(x)\big].$
We compute $U(D)$ separately for the code-mixed condition ($\mathrm{CM}$) and the translated condition ($T(\mathrm{CM})$), using the same prompts and decoding settings as in Section 9. Averaged over all ten languages on the Local TestSet, we observe that the difference in utility between the two conditions is small:
$U_{\mathrm{CM}}=0.87,\qquad U_{T(\mathrm{CM})}=0.86,$
with all per-language changes within $\pm 2.5$ percentage points. In other words, the translation intervention does not systematically degrade the model’s ability to produce direct, on-topic answers, even though it substantially reduces Attack Success Rates. This supports our claim in Section 9 that the intervention primarily suppresses unsafe generations rather than trivially “shutting down” the model.
Appendix C Sensitivity analysis of mixing ratios
While our primary evaluation utilizes the Matrix Language Frame (MLF) with a 60:40 ratio to adhere to established sociolinguistic constraints for grammatical validity, real-world web interactions particularly on social media that exhibit fluid code-mixing behaviors do not adhere to such strict ratios. To verify that our findings of attributional collapse are robust and not artifacts of a specific syntactic configuration, we conducted a sensitivity analysis varying the mixing intensity.
We evaluate safety degradation across a spectrum of mixing ratios, ranging from 80:20 (Matrix dominant) to 20:80 (Embedded dominant) using the Bengali-English dataset. The results, summarized in Table 7, reveal a strong monotonic relationship between the degree of linguistic perturbation and the severity of safety failure.
Ratio phi-4B (%) mistral-7B (%) llama-3.1-8B (%)
80:20 59.2 66.8 48.1
70:30 63.7 69.9 52.3
50:50 69.0 76.4 64.7
20:80 75.2 81.6 70.9
Table 7. ASR sensitivity to mixing ratios. The table demonstrates a consistent degradation in safety as the complexity of the code-mixing increases (moving away from the Matrix language dominance).
C.1. Implications for web safety
This monotonic degradation confirms that attributional safety failures are distributional, not incidental. The 60:40 ratio chosen for the main body of this work serves as a conservative, representative baseline. The fact that safety deteriorates even further at higher mixing ratios (e.g., 20:80) suggests that our main results may actually underestimate the risk in highly informal, chaotic web environments where code-mixing ratios fluctuate wildly. This underscores the necessity for defenses that are robust to varying degrees of linguistic interference, rather than those tuned to specific syntactic templates.
Appendix D Cultures and categories
Figure 12. Utilized cultures and key areas that could lead to potential harm within each culture.
(a) Loss of attribution (code-mixed Bengali)
(b) Loss of attribution (code-mixed Hindi)
Figure 13. Case 1: English inputs yield non-harmful responses, while Bengali and Hindi code-mixed variants triggered harmful outputs. Both word clouds show harmful tokens (e.g., disgusting, parasites) that lose attribution under code-mixing.
(a) Unaltered attribution (code-mixed Bengali)
(b) Unaltered attribution (code-mixed Hindi)
Figure 14. Case 2: Both English and code-mixed inputs produce harmful responses. The word clouds show that harmful tokens retain high attribution, indicating stable but unsafe behavior across modalities.
(a) Unaltered attribution (code-mixed Bengali)
(b) Unaltered attribution (code-mixed Hindi)
Figure 15. Case 3: Both English and code-mixed inputs result in non-harmful responses. The word clouds show minimal attribution drift, indicating stable and safe behavior across both forms.
Additional analyzes and diagnostics
This appendix provides focused evaluations to further support the arguments in the main text.
D.1. Token drift visualization across cultures (Case 1)